PROSPECTIVE WORK FOR ALMA: THE MILLIMETERWAVE AND SUBMILLIMETERWAVE SPEC-TRUM OF ^{13}C-GLYCOLALDEHYDE

IMANE HAYKAL, LAURENT MARGULÈS, THERESE R. HUET, ROMAN MOTIYENKO, Laboratoire PhLAM, UMR8523 CNRS-Université Lille 1, F-59655 Villeneuve d’Ascq Cedex, France; and J.-C. GUILLEMIN, UMR6226 CNRS-Ecole Nationale Supérieure de Chimie de Rennes,F-35700 Rennes, France.

Glycolaldehyde has been identified in interstellar sources^{a}. The relative abundance ratios of the three isomers (acetic acid) : (glycolaldehyde) : (methylformate) were estimated . The detection of ^{13}C_{1} and ^{13}C_{2} isotopomers of methylformate has been recently reported in Orion, as a result of the detailled labororatory spectroscopic study^{b}. Therefore the spectroscopy of the ^{13}C isotopomers of glycolaldehyde is investigated in laboratory in order to provide data for an astronomical search. The instrument ALMA will certainly be a good instrument to detect them. Up to now, only the microwave spectra of 13CH_{2}OH-CHO and of CH_{2}OH-^{13}CHO have been observed several years ago in the 12-40 GHz range^{c}. Spectra of both species are presently recorded in Lille in the 150-950 GHz range with the new submillimetre-wave spectrometer based on harmonic generation of a microwave synthesizer source, using only solid-state devices, and coupled to a cell of 2.2 m length^{d}. The absolute accuracy of the line positions is better than 30 KHz. The rotational structure of the ground state and of the three first excited vibrational states has been observed. Two ^{13}C enriched samples were used. The analysis is in progress. This work is supported by the Programme National de Physico-Chimie du Milieu Interstellaire (PCMI-CNRS) and by the contract ANR-08-BLAN-0054

aJ. M. Hollis, S. N. Vogel, L. E. Snyder, et al., Astrophys. J. 554(2001) L81 ; R. A. H. Butler, F. C. De Lucia, D. T Petkie, et al., Astrophys. J. Supp. 134 (2001) 319 ; M. T. Beltran, C. Codella, S. Viti, R. Niri, R. Cesaroni, Astrophys. J. 690 (2009) L93.

bM. Carjaval, L. Margulès, B. Tercero et al., Astron. Astrophys. 500 (2009) 1109.

cK.-M. Marstokk and H. Møllendal, J. Mol. Struct. 16 (1973) 259.

dR. A. Motiyenko, L. Margulès, E. A. Alekseev et al., J. Mol. Spectrosc. 264 (2010) 94.